Reinegar Corporation is planning two new issues of 25-year bonds. Bond Par will be sold at its $1,000 par value, and it will have a 10% semiannual coupon. Bond OID will be an Original Issue Discount bond, and it will also have a 25-year maturity and a $1,000 par value, but its semiannual coupon will be only 6.25%. If both bonds are to provide investors with the same effective yield, how many of the OID bonds must Reinegar issue to raise $3,000,000? Disregard flotation costs, and round your final answer up to a whole number of bonds.

a.

4,337

b.

4,562

c.

4,676

d.

4,228

e.

4,448
Answer #1

For Bond Par bond, since the bond is selling at PAR, its YTM is equal to coupon rate

So YTM for both the bond = 10% annual, compounded semiannually.

Given for OID Bond

Face value = $1000

coupon rate = 6.25% semiannual

So coupon = 6.25%/2 of 1000 = $31.25

YTM = 10%

Years to maturity = 25 years,

Using financial calculator to solve for Price of the bond,

FV =1000

PMT = 31.25

N = 25*2 = 50

I/Y = 10/2 = 5

compute PV, we get PV = -657.70

So bond will sel at a price of $657.70

since company has to raise 3000000, they need to sell 3000000/657.7 = 4561.34 bonds

So total bond sold = 4562

So, option (b) is correct.